A 77.18 g sample of a hydrocarbon (CxHy) has a mass of carbon that is equal to 66.89 g. Given this information, determine the empirical formula of this hydrocarbon. MMC = 12.01 and MMH = 1.008) 

Throughout your calculations, maintain at least 4 significant figures. 


In each answer box place the appropriate subscript for each spoiled element which in essence conveys this compound's empirical formula. 
Answer #1

Mass of carbon =66.89 g

Mass of Hydrogen = 77.18-66.89 =10.29 g

Moles of carbon, x=66.89/12.01 = 5.569

Moles of Hydrogen, y=10.29/1.008 = 10.21

C5.569 H10.21

Empirical formula is C1H1.833

C 1 H 1.833